OpenStudy (anonymous):

What is the slope of the line that contains the points (-6, -7) and (6, -5)? Please help

OpenStudy (igreen):

Plug the two points in the slope formula. \(m = \dfrac{y_2-y_1}{x_2-x_1}\) \(m = \dfrac{-5+7}{6+6}\) Add: \(m = \dfrac{2}{12}\) Can you simplify \(\dfrac{2}{12}\)?

OpenStudy (igreen):

@anubisfang

OpenStudy (anonymous):

yes its 1/6

OpenStudy (igreen):

Yep, you got it..so the slope is 1/6.

OpenStudy (igreen):

Welcome to Open Study! You can give medals by clicking 'Best Response'.

OpenStudy (anonymous):

thank you